OpenStudy (anonymous):

If (a/b) = (c/d), which of these statements is not true? ac = bd, given that b and d are not 0. b/a = d/c, given that a, b, c, and d are not 0. ad = bc, given that b and d are not 0. (a + b)/b = (c + d)/d, given that b and d are not 0.

OpenStudy (antonio_xx2):

explain this plese

OpenStudy (anonymous):

well the first sentence is a question the rest are answers

OpenStudy (antonio_xx2):

ok

OpenStudy (anonymous):

these are the answers ac = bd, given that b and d are not 0. b/a = d/c, given that a, b, c, and d are not 0. ad = bc, given that b and d are not 0. (a + b)/b = (c + d)/d, given that b and d are not 0.

OpenStudy (antonio_xx2):

ac=bd t because a/b=c/d then a=(cb)/d and ac=(c^2b)/d which is not bd because when you substitute, bd=(d^2a)/c, unless d=c and a=b

OpenStudy (anonymous):

thank you <3

OpenStudy (antonio_xx2):

sure